OpenStudy (anonymous):

Someone check my answers please? Multiply or find the special product: 1. (7x-2)(4x-3) my answer: 28x^2-29x+6 2. (4a+5b)(4a-5) m answer: 16a^2-25b^2 3. (x^2-x+1)(x^2-3x-2) my answer: x^4-4x^3+2x^2-x-2 4. (5+sqrt x)(5-sqrt x) my answer: 25-x Thank you!

OpenStudy (uri):

(7x-2)(4x-3) 7x(4x-3)-2(4x-3) 28x^2-21x-8x+6 28x^2-29x+6 Correct!

OpenStudy (anonymous):

Thank you!! Do you know if the other ones are right as well??

OpenStudy (uri):

(4a+5b)(4a-5) 4a(4a-5)5b(4a-5) 16a^2-20a+20ab-25b

OpenStudy (anonymous):

aww...I knew there had to be a typo somewhere! #2 is supposed to be (4a+5b)(4a-5b)

OpenStudy (anonymous):

I just checked the rest of them and i'm pretty sure they are all right! Yay! Thank you for your help! :)

OpenStudy (austinl):

\(\huge \color{red}{\text{FOIL}}\) \(\color{red}{F}irst\) \(\color{red}{O}uter\) \(\color{red}{I}nner\) \(\color{red}{L}ast\) \((ax+b)(cx+d)=acx^2+adx+bcx+bd\) \((7x-2)(4x-3)=(7x\times4x)+(3\times7x)+((-2)\times4x)+((-2)\times(-3))\) \(=28x^2-29x+6\) As long as you have down that basic form you should be alright. :)